Suppose you observe the following situation:

  Security

Beta

Expected Return

  Pete Corp.

1.30

.140

  Repete Co.

.99

.113

  1. Assume these securities are correctly priced. Based on the CAPM, what is the expected return on the market? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
  2. What is the risk-free rate? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

Expected return=risk free rate+beta*(market rate-risk free rate)

14=Rf+1.3*(Rm-Rf)

14=1.3*Rm-0.3Rf

Rm=(14+0.3Rf)/1.3

Also:

11.3=Rf+0.99*(Rm-Rf)

11.3=0.99Rm+0.01Rf

11.3=0.99*(14+0.3Rf)/1.3+0.01Rf

11.3=10.6615385+0.228461538Rf+0.01Rf

Rf=(11.3-10.6615385)/(0.228461538+0.01)

=2.68%(Approx)=risk free rate

Rm=(14+0.3Rf)/1.3

=11.39%(Approx)=market rate